# Learning the Language of Protein Structure

### v1

@@ 1 Introduction
Currently, the deep learning landscape is increasingly converging towards a unified paradigm centered around attention-based architectures (Vaswani et al., 2017) and sequence modeling. This shift has been driven by the impressive performance and scalability of transformers, and even accelerated since treating non-standard data modalities as sequence-modeling problems has proven highly effective. Indeed, transformers-based models are leading methods in many machine learning domains, including image representation (Radford et al., 2021), image (Chen et al., 2020; Chang et al., 2022) and audio generation (Ziv et al., 2024), and for Reinforcement Learning (Chen et al., 2021; Boige et al., 2023). This trend has greatly benefited sequence-based biological models, allowing for the direct application of NLP methodologies like GPT (Radford and Narasimhan, 2018) and BERT (Devlin et al., 2019) with notable success (Ferruz et al., 2022; Lin et al., 2023).
@@
In particular, large multi-modal models (LMMs), leveraging transformer backbones, are emerging as a key tool with applications in various fields such as: ubiquitous AI (GPT4 (Achiam et al., 2023), LLaVA (Liu et al., 2024), Gemini (Gemini et al., 2023), Flamingo (Alayrac et al., 2022)); text-conditioned generation of images (Parti (Yu et al., 2022b), Muse (Chang et al., 2023)) or sounds (MusicGen (Copet et al., 2023)); and even reinforcement learning (Reed et al., 2022). LMMs are also instantiated in biological settings, such as medicine (MedLlama (Xie et al., 2024), Med-Gemini (Yang et al., 2024b) Med-PaLM (Tu et al., 2024)) and genomics (ChatNT (Richard et al., 2024)). Core to all these LMMs is the use of pre-trained encoders as a mechanism for combining modalities in sequence space. For instance, LLaVa (Liu et al., 2024) and Flamingo (Alayrac et al., 2022) used pre-trained vision encoders, ViT-L/14 (Radford et al., 2021) and Normalizer Free ResNet (NFNet) (Alayrac et al., 2020) respectively, while Copet et al. (2023) leverages the the codec of Défossez et al. (2023). In general, training robust representations of specific modalities facilitates their incorporation into state-of-the-art LMMs. However, to the best of our knowledge, there is no established methodology that readily allows the application of sequence-modelling to protein structures.
@@
Foldseek (van Kempen et al., 2024) proposes a quantized autoencoder to encode local protein geometry which proved successful for database search and to some extent, in LMMs by augmenting the residue-based vocabulary (Su et al., 2024). However, since it only extracts local features at a residue level, it cannot provide global representation of the protein structures, preventing its use for instance for structure generation, or binding prediction for which including global information is necessary (Krapp et al., 2023). Despite these advances in related domains, structure-based modeling of biological data such as proteins remains a formidable challenge. Unlike sequences, protein structures are inherently three-dimensional and continuous, which complicates the direct application of transformer models that primarily handle discrete data. Instead, structure-based methods often design bespoke geometric deep learning methodologies to process Euclidean data; for example graph-neutral network encoders (Dauparas et al., 2022; Krapp et al., 2023) and structurally-aware modules such as those in AlphaFold (Jumper et al., 2021). Moreover, generative modelling of structures is typically performed with methods designed for continuous variables; such as diffusion (Watson et al., 2023; Yim et al., 2023) and flow matching (Bose et al., 2024), rather than the discrete-variable models that have proved so successful in sequence modelling.
@@
Our contributions are threefold. First, we introduce a series of quantized autoencoders that effectively discretize protein structures into sequences of tokens while preserving the necessary information for accurate reconstruction. Second, we validate our autoencoders through qualitative and quantitative analysis, and various ablation studies, supporting our design choices. Third, we demonstrate the efficacy and practicality of the learned representations with experimental results from a simple GPT model trained on our learned codebook, which successfully generates novel, diverse, and structurally viable protein structures. All source code for experiments is publicly available online [^1] Redacted to maintain anonymity.
@@ 2.1 Protein Structure Autoencoder
The task our decoder addresses formulates more broadly as a sequence to point-cloud task. A paradigmatic example of such a task in biology is the protein folding problem, where the confirmation of a protein is estimated given its primary sequence of amino-acids. Jumper et al. (2021) successfully tackled this task proposing a novel architecture for point cloud estimation from latent sequence of embeddings.
@@ 2.2 Training
We use approximately 310000 entries available in the Protein Data Bank (PDB) (Berman et al., 2000) as training data. The presence of large groups of proteins causes imbalances in the data set, with many protein from the same family sharing structural similarities. To mitigate this issue, we sample the data inversely proportional to the size of the cluster it belongs to when clustering the data by sequence similarity using MMseqs2222The cluster size is readily available in the PDB data set: https://www.rcsb.org/docs/grouping-structures/sequence-based-clustering. We filter all chains shorter than 50 residues and crop the structures that have more than 512 residues by randomly selecting 512 consecutive amino acids in the corresponding sequences. We randomly select $90\%$ for training and use the remaining as test set. Amongst these $10\%$ withhold protein structures, we retain $20\%$ for validation, the remaining $80\%$ being used for test.
@@ Results
Downsampling Ratio Number of Codes Compression Factor Results
RMSD ($\downarrow$) TM-Score ($\uparrow$)
1 4096 48 1.89 Å 0.93
64000 48 1.59 Å 0.95
without quantization -- 0.97 Å 0.98
2 4096 96 2.79 Å 0.87
64000 96 2.31 Å 0.91
without quantization -- 1.45 Å 0.95
4 4096 144 4.63 Å 0.79
64000 144 3.23 Å 0.85
without quantization -- 2.19 Å 0.91
Table 1: Test set reconstruction results of our discrete auto-encoding method for several down-sampling ratio and (implicit) codebook size. Note that a RMSD below $2$Å  is considered of the order of experimental resolution and two proteins with a TM-score $>0.5$ are considered to have the same fold. The compression is defined as: the number of bits necessary to store the $N\times 4\times 3$ backbone positions divided by the number of bits necessary to store the $\frac{N}{r}$ tokens
Table 1 clearly indicates that increasing the downsampling factor or decreasing the codebook size correspondingly impacts the reconstruction accuracy. This is expected as it essentially enforces greater compression in our autoencoder leading to a loss of information - nevertheless in all cases we find that the achievable reconstruction performance is still within a few angstrom with TM-scores clearly exceeding the 0.5 threshold on average. The fact that the performances improve with increasing the codebook size also shows that our method does not suffer from codebook collapse (i.e. only a subset of the codes being used by the trained model), an issue well known and documented with other quantization methods (Huh et al., 2023; Takida et al., 2024). Finally, comparing to continuous autoencoders, our learned quantization demonstrates significant information compression at the expense of only a small decrease in the reconstruction precision of the order of $0.5-1$.
@@ 4 Related Works
("Eguchi et al., 2022) learns a VAE using as inputs matrix distances and predicting 3D coordinates. The supervision is done by comparing distance matrices. Despite the straightforward nature of sampling from a VAE latent space, it often yields subpar sample quality, see (Kingma et al., 2016). Foldseek (van Kempen et al., 2024) propose to learn a VQ-VAE on small chunks of protein comparing true and reconstructed local geometrical features. Because of the locality of the data used in Foldseek, decoding only small chunks of structures, this quantization defines an interesting tool for efficient structure database search but does not enable full backbone structure sampling. Closer to our work, Gao et al. (2023) propose a quantized autoencoder of protein structures, with however only limited reconstruction performances limiting its applicability.
@@
A substantial body of literature addresses the challenge of sampling the protein structure space. Numerous studies have advocated for the use of diffusion-based models Wu et al. (2024); Yim et al. (2023); Watson et al. (2023) or flow matching techniques (Bose et al., 2024). While many of these works, such as those by (Yim et al., 2023; Watson et al., 2023; Bose et al., 2024), employ complex architectures to ensure invariance to rigid-body transformations, (Wu et al., 2024) opted for an alternative parameterization directly preserving symmetries allowing the authors to capitalize on conventional architectures, yet working only on small protein crops.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
This research was supported with Cloud TPUs from Google’s TPU Research Cloud (TRC).
@@ A.1 Architectures
Algorithm 2 Resampling Layer
@@ A.2 Training Details and Metrics
A.3 De novo Structure Generation
A.3.1 Training details
@@
A.3.2 Structure generation metrics

### v2

@@ 1 Introduction
Currently, the deep learning landscape is increasingly converging towards a unified paradigm centered around attention-based architectures (Vaswani et al., 2017) and sequence modeling. This shift has been driven by the impressive performance and scalability of transformers, and even accelerated since treating non-standard data modalities as sequence-modeling problems has proven highly effective. Indeed, transformers-based models are leading methods in many machine learning domains, including image representation (Radford et al., 2021), image (Chen et al., 2020; Chang et al., 2022) and audio generation (Ziv et al., 2024), and for Reinforcement Learning (Chen et al., 2021; Boige et al., 2023). This trend has greatly benefited sequence-based biological models, allowing for the direct application of NLP methodologies like GPT (Radford and Narasimhan, 2018) and BERT (Devlin et al., 2019) with notable success (Ferruz et al., 2022; Lin et al., 2023b).
@@
In particular, large multi-modal models (LMMs), leveraging transformer backbones, are emerging as a key tool with applications in various fields such as: ubiquitous AI (GPT4 (Achiam et al., 2023), LLaVA (Liu et al., 2024b), Gemini (Gemini et al., 2023), Flamingo (Alayrac et al., 2022)); text-conditioned generation of images (Parti (Yu et al., 2022b), Muse (Chang et al., 2023)) or sounds (MusicGen (Copet et al., 2023)); and even reinforcement learning (Reed et al., 2022). LMMs are also instantiated in biological settings, such as medicine (MedLlama (Xie et al., 2024), Med-Gemini (Yang et al., 2024b) Med-PaLM (Tu et al., 2024)) and genomics (ChatNT (Richard et al., 2024)). Core to all these LMMs is the use of pre-trained encoders as a mechanism for combining modalities in sequence space. For instance, LLaVa (Liu et al., 2024b) and Flamingo (Alayrac et al., 2022) used pre-trained vision encoders, ViT-L/14 (Radford et al., 2021) and Normalizer Free ResNet (NFNet) (Alayrac et al., 2020) respectively, while Copet et al. (2023) leverages the the codec of Défossez et al. (2023). In general, training robust representations of specific modalities facilitates their incorporation into state-of-the-art LMMs. However, to the best of our knowledge, there is no established methodology that readily allows the application of sequence-modelling to protein structures.
@@
Despite these advances in related domains, structure-based modeling of biological data such as proteins remains a formidable challenge. Unlike sequences, protein structures are inherently three-dimensional and continuous, which complicates the direct application of transformer models that primarily handle discrete data. Instead, structure-based methods often design bespoke geometric deep learning methodologies to process Euclidean data; for example graph-neutral network encoders (Dauparas et al., 2022; Krapp et al., 2023) and structurally-aware modules such as those in AlphaFold (Jumper et al., 2021). Moreover, generative modelling of structures is typically performed with methods designed for continuous variables; such as diffusion (Watson et al., 2023; Yim et al., 2023) and flow matching (Bose et al., 2024), rather than the discrete-variable models that have proved so successful in sequence modelling.
@@
Our contributions are threefold. First, we introduce a series of quantized autoencoders that effectively discretize protein structures into sequences of tokens while preserving the necessary information for accurate reconstruction. Second, we validate our autoencoders through qualitative and quantitative analysis, and various ablation studies, supporting our design choices. Third, we demonstrate the efficacy and practicality of the learned representations with experimental results from a simple GPT model trained on our learned codebook, which successfully generates novel, diverse, and structurally viable protein structures. All source code for experiments is publicly available online at: https://github.com/instadeepai/protein-structure-tokenizer/.
@@ 2.1 Protein Structure Autoencoder
The task our decoder addresses formulates more broadly as a sequence to point-cloud task. A paradigmatic example of such a task in biology is the protein folding problem, where the confirmation of a protein is estimated given its primary sequence of amino-acids. Jumper et al. (2021) successfully tackled this task proposing a novel architecture for point cloud estimation from latent sequence of embeddings. Therefore, we use AlphaFold-2 structure module to parameterize our decoder and learning the parameter from scratch.
@@ 2.2 Training
We use approximately 310000 entries available in the Protein Data Bank (PDB) (Berman et al., 2000) as training data. The presence of large groups of proteins causes imbalances in the data set, with many protein from the same family sharing structural similarities. To mitigate this issue, we sample the data inversely proportional to the size of the cluster it belongs to when clustering the data by sequence similarity using MMseqs2111The cluster size is readily available in the PDB data set: https://www.rcsb.org/docs/grouping-structures/sequence-based-clustering. We filter all chains shorter than 50 residues and crop the structures that have more than 512 residues by randomly selecting 512 consecutive amino acids in the corresponding sequences. We randomly select $90\%$ of the clusters for training and use the remaining as test set. Amongst these $10\%$ withhold protein structures clusters, we retain $20\%$ for validation, the remaining $80\%$ being used for test.
@@ Results
Table 1: Average Test set reconstruction results of our discrete auto-encoding method for several down-sampling ratios and (implicit) codebook sizes. For CASP-15 we report the median of the metrics due to the limited dataset size. Note that a RMSD below $2$Å  is considered of the order of experimental resolution and two proteins with a TM-score $>0.5$ are considered to have the same fold. The compression is defined as: the number of bits necessary to store the $N\times 4\times 3$ backbone positions divided by the number of bits necessary to store the $\frac{N}{r}$ tokens multiplied by $\log_{2}(\#\,Codes)$
Downsampling Ratio Number of Codes Compression Factor Results CASP15 RMSD ($\downarrow$) TM-Score ($\uparrow$) RMSD ($\downarrow$) TM-Score ($\uparrow$) 1 432 88 2.09 Å 0.91 1.75 Å 0.89 1728 71 1.79 Å 0.93 1.33 Å 0.94 4096 64 1.55 Å 0.94 1.25 Å 0.94 64000 48 1.22 Å 0.96 0.94 Å 0.97 without quantization -- 0.97 Å 0.98 1.07 Å 0.93 2 4096 128 2.22 Å 0.90 1.73 Å 0.89 64000 96 1.95 Å 0.92 1.82 Å 0.90 without quantization -- 1.45 Å 0.95 1.44 Å 0.90 4 4096 256 4.10 Å 0.81 2.79 Å 0.77 64000 192 2.96 Å 0.86 2.55 Å 0.80 without quantization -- 2.19 Å 0.91 1.98 Å 0.84
Table 1 clearly indicates that increasing the downsampling factor or decreasing the codebook size correspondingly impacts the reconstruction accuracy. This is expected as it essentially enforces greater compression in our autoencoder leading to a loss of information - nevertheless in all cases we find that the achievable reconstruction performance is still within a few angstrom with TM-scores clearly exceeding the 0.5 threshold on average. The fact that the performances improve with increasing the codebook size also shows that our method does not suffer from codebook collapse (i.e. only a subset of the codes being used by the trained model), an issue well known and documented with other quantization methods (Huh et al., 2023; Takida et al., 2024). This is also noticeable in Figure 2 that shows that, given a downsampling ratio, larger codebooks effectively decrease the reconstruction errors. Finally, comparing to continuous autoencoders, our learned quantization demonstrates significant information compression at the expense of only a small decrease in the reconstruction precision of the order of $0.5-1$. Additionally, we provide in Section A.3 detailed distribution of the RMSD and TM-Scores for both CASP-15 and the held-out test set. Notably, we can see that increasing the downsampling ratio, or decreasing the codebook size tends to thickens the right tail (resp. the left) of the distribution for the RMSD (resp. the TM score).
Figure 2: Evolution of the RMSD (left) and TM-score (right) distribution with the codebook size for a dowsampling ratio of 1 on CASP-15 data.
@@ 4 Related Works
("Eguchi et al., 2022) learns a VAE using as inputs matrix distances and predicting 3D coordinates. The supervision is done by comparing distance matrices. Despite the straightforward nature of sampling from a VAE latent space, it often yields subpar sample quality, see (Kingma et al., 2016).
Discrete representation learning for protein structures has recently garnered increasing attention. Foldseek (van Kempen et al., 2024) introduces a quantized autoencoder to encode local protein geometry, demonstrating success in database search tasks. However, as it focuses solely on local features at the residue level, it lacks the capacity to provide global representation of protein structures. This limitation restricts its application in tasks like structure generation or binding prediction, where global information is critical (Krapp et al., 2023). Building on the 3Di-alphabet introduced by Foldseek, Su et al. (2024); Heinzinger et al. (2023) propose structure-aware protein language models that integrate structure tokens with sequence tokens. Additionally, Li et al. (2024) combines a structural autoencoder with K-means clustering applied to the latent representation of a fixed reference dataset.
More closely related to our work, Gao et al. (2023) adapts VQ-VAE (van den Oord et al., 2017) for protein structures. However, its limited reconstruction performance constrains its applicability. Lin et al. (2023a) explores discrete structural representation learned by such VQ-VAEs (van den Oord et al., 2017), while Liu et al. (2023) trains a diffusion model on the discrete latent space derived from this approach. Similarly, and concurrently with our work, Liu et al. (2024a) combines finite scalar quantization (Mentzer et al., 2024) with a specialized transformer-based autoencoder for proteins, RNA, and small molecules.
Very recently, efforts have emerged to combine quantized structural representation with discrete sequence representation, enabling multimodal generative models trained on a joint discrete latent space. FoldToken (Gao et al., 2024b; a), is a concurrent approach that shares conceptual similarities with our work but differs in key methodological and application-oriented aspects. FoldToken employs joint quantization of sequence and structure, enabling integration across modalities, whereas our method focuses exclusively on structural information. This decoupling allows for mode-specific pretraining, aligning with strategies from subsequent works (Hayes et al., 2024; Lu et al., 2024). Methodologically, FoldToken introduces a series of improvements to existing VQ methods (van den Oord et al., 2017) aimed at enhancing reconstruction accuracy; whereas we adopt the FSQ framework, which reduces the straight-through gradient estimation gap inherent to VQ methods while improving codebook utilization. Furthermore, while FoldToken primarily emphasizes backbone inpainting and antibody design (Gao et al., 2024b; a), our work considers de novo generation of complete structures.
@@
A substantial body of literature addresses the challenge of sampling the protein structure space. Numerous studies have advocated for the use of diffusion-based models Wu et al. (2024); Yim et al. (2023); Watson et al. (2023) or flow matching techniques (Bose et al., 2024). While many of these works, such as those by (Yim et al., 2023; Watson et al., 2023; Bose et al., 2024), employ complex architectures to ensure invariance to rigid-body transformations, (Wu et al., 2024) opted for an alternative parameterization directly preserving symmetries allowing the authors to capitalize on conventional architectures, yet working only on small protein crops. Recently, Wang et al. (2024b) employs a lookup-free quantizer (Yu et al., 2024) as a structure tokenizer and trains a diffusion protein language model (Wang et al., 2024a) on the concatenated sequence and structure tokens. Other works (Hayes et al., 2024; Lu et al., 2024) simply uses VQ-VAEs (van den Oord et al., 2017) to tokenize the structures and train large language models (LLM) on the combination of sequence and structure tokens.
@@ A.1 Architectures
Algorithm 2 Resampling Layer with Positional Encoding
1:Input: $\texttt{Inputs}\in\mathbb{R}^{T\times d}$ , Mask, target size: p, features dim: d , Queries: [Optional]
[…]
@@ A.2 Training Details and Metrics
A.3 Additional Results: Structures Autoencoding
Figure 6: Evolution of the RMSD distribution on CASP-15 dataset with the downsampling ratio, with a fixed codebook size of 4096
Figure 7: Evolution of the TM-Score distribution on CASP-15 dataset with the downsampling ratio, with a fixed codebook size of 4096
Figure 8: Evolution of the RMSD distribution on CASP-15 dataset with the downsampling ratio, with a fixed codebook size of 4096
Figure 9: Evolution of the TM-Score distribution on CASP-15 dataset with the downsampling ratio, with a fixed codebook size of 4096
Figure 10: RMSD (left) and TM-Score (right) distribution on the held out test set for the codebook size of 432 and downsampling ratio of 1.
Figure 11: RMSD (left) and TM-Score (right) distribution on the held out test set for the codebook size of 1728 and downsampling ratio of 1.
Figure 12: RMSD (left) and TM-Score (right) distribution on the held out test set for the codebook size of 4096 and downsampling ratio of 1.
Figure 13: RMSD (left) and TM-Score (right) distribution on the held out test set for the codebook size of 64,000 and downsampling ratio of 1.
Figure 14: RMSD (left) and TM-Score (right) distribution on the held out test set for the codebook size of 4096 and downsampling ratio of 2.
Figure 15: RMSD (left) and TM-Score (right) distribution on the held out test set for the codebook size of 64,000 and downsampling ratio of 2.
Figure 16: RMSD (left) and TM-Score (right) distribution on the held out test set for the codebook size of 4096 and downsampling ratio of 4.
Figure 17: RMSD (left) and TM-Score (right) distribution on the held out test set for the codebook size of 64,000 and downsampling ratio of 4.
A.4 De novo Structure Generation
A.4.1 Training details
@@
A.4.2 Structure generation metrics
@@ A.4.4 Additional Results
Figure 22: Evolution with the length of the generated sequence of the per-residue negative Log-Likelihood of the selected amino acids as provided by ProteinMPNN